$4.52/gal to ¢/mL


with the correct number of sig figs and units


1
Expert's answer
2021-03-01T02:35:25-0500

Q167566

Convert $4.52/gal to ¢/mL

with the correct number of sig figs and units.


Solution:


For converting $4.52/gal to ¢/mL, we will require two conversion factors.

One for converting '$' to '¢' and

second for converting 'gal' to 'mL'.


1 dollar = 100 cent and 1 gallon = 3785 mL.


Using both this conversion factor we can convert $4.52/gal to ¢/mL.


"cent/mL = \frac{4.52 dollar }{1 gallon} * \frac{100 cent }{ 1 dollar} * \frac{1 gallon }{3785 mL } = 0.119 cent /mL"


which can also be written as 0.119 ¢/mL.





In the question, we are given $4.52/gal in 3 significant figures.

So our final answer must also be in 3 significant figures.


So the final answer is 0.119 ¢/mL.


----------------x -------------------------------x--------------------------------------------------x--------------


We can also use





Thank you. :)